# D-LGP: Dynamic Logic-Geometric Program for Combined Task and Motion Planning

### v1

@@ III-C D-LGP Algorithm
To tackle (1) iteratively from the target configuration, we introduce Dynamic Logic-Geometric Program (D-LGP) in Alg. 2. Here, we employ a pseudoinverse successor function, denoted as $succ^{\dagger}(s_{k},s_{k+1})$, to deduce the goal-oriented action $a_{k+1}$ for transitioning from $s_{k}$ to $s_{k+1}$. This function disregards the symbolic and geometric constraints, focusing solely on determining the necessary action to reach the target state $s_{k+1}$ from $s_{k}$.
Given the current configuration $\bm{x}((k-1)T)$, we employ the conflict-driven TaskGraph algorithm (Alg. 1) to assess the feasibility of the proposed action. If the action is found infeasible, we generate a sequence of actions ($\mathcal{B}_{a}$) and subgoals ($\mathcal{B}_{g}$) to construct a tree connecting $s_{k-1}$ to $s_{K}$. $sub\_goal(\cdot)$ identifies the crucial subgoals that must be accomplished first to render action $a_{k}$ feasible, given the current state $\bm{x}((k-1)T)$. For example, if the final target is to grasp object A within a cluttered environment flanked by taller blocks B and D, grasping B and D emerges as essential subgoals. It is important to note that $\mathcal{B}_{g}$ does not encompass all the necessary subgoals between $s_{k-1}$ and $s_{K}$. $\mathcal{B}_{g}$ and $\mathcal{B}_{a}$ only guarantee the feasibility of the last subgoal and action with respect to the current configuration. Subsequently, we calculate the continuous path from $t=(k-1)T$ to $t=kT$ based on (4). Since computing (4) is expensive, we decide to only focus on the final pose at $t=kT$ for fast feasibility checking and cost computation. Based on this, we can quickly solve it and output the placement locations $\bm{u}(kT)$ or returns global infeasibility.
This process is run repeatedly until the full action skeleton $\{a_{1},\dots,a_{K}\}$ connecting $s_{0}$ and $s_{K}$ are found. A full path optimization will then be solved to find the global optimal trajectory, focusing on the keyframes $t=0,T,\dots,KT$. This idea coincides with the pose bound $\mathcal{P}_{pose}$ and sequence bound $\mathcal{P}_{seq}$ used in multi-bound tree search [3], which proves using coarse path discretization can effectively speed up the optimization process. The difference is that we eliminate the full path bound $\mathcal{P}_{full}$ to fully utilize the robot controller to cope with uncertainties and external disturbance in the real world. Our TAMP approach runs in Cartesian space and relies on Operational Space Controller to actuate the robot joints. This is also highlighted in [18], where a Cartesian formulation of LGP was proposed to achieve object-centric TAMP.
@@
2:$\bm{x}(kT):$ current geometric state, $s_{k}:$ current symbolic state
@@
5:$a_{k+1}=succ^{\dagger}(s_{k},s_{K})$
6:$\mathcal{B}_{a}=[a_{k+1}],\mathcal{B}_{g}=[g_{K}]$
7:while $(\bm{x}(kT),a_{k+1})$ infeasible do
8:   $g^{*}=sub\_goal(\bm{x}(kT),a_{k+1})$triangleright subgoal reasoning
9:   $s^{*}\in\mathcal{G}(g^{*})$
10:   $a_{k+1}=succ^{\dagger}(s_{k},s^{*})$
11:   $\mathcal{B}_{a}.append(a_{k+1})$
12:   $\mathcal{B}_{g}.append(g^{*})$
13:end while
@@
1:         Level 1: Dynamic Tree Search
2:Initialize:
@@
4:$\mathcal{B}_{a},\mathcal{B}_{g}\leftarrow\text{TaskGraph}(\bm{x}_{0},s_{0})$
5:while not solved do
6:   $a_{k}=\mathcal{B}_{a}[-1]$ (feasible action given $\bm{x}_{k-1}$)
7:   $s_{k}=succ(s_{k-1},a_{k})$
8:   $g_{k}=\mathcal{B}_{g}[-1]$
9:   $\bm{x}(kT),\bm{u}(kT)=\text{MIQP}(a_{k},g_{k})$triangleright Eq. (4)
10:   $\mathcal{B}_{a}^{{}^{\prime}},\mathcal{B}_{g}^{{}^{\prime}}=\text{TaskGraph}(\bm{x}(kT),s_{k})$triangleright see Alg. 1
11:   $\mathcal{B}_{a}\leftarrow\mathcal{B}_{a}^{{}^{\prime}}$, $\mathcal{B}_{g}\leftarrow\mathcal{B}_{g}^{{}^{\prime}}$
12:   $k\leftarrow k+1$
13:end while
14:Level 2: Full Optimization
15:Given $a_{1:K}$:
16:
@@
18:
@@ IV-A Benchmarks
We test our method in 3 different tabletop domains, all requiring long-horizon physical reasoning about the discrete modes and continuous motions. The reward is only given at the termination of the whole process. Across all task domains, we employ a random selection of 15 distinct initializations to assess the performance of our algorithm.
a) Obstructed Pick X (OP-X): In a cluttered environment, the robot’s task is to grasp an object surrounded by X-1 taller objects. If the taller objects are located near the target object, the robot must relocate them to create a clear path for grasping. The reward is solely tied to the final configuration, specifically whether the target object is successfully grasped with no collisions. This necessitates the robot’s comprehensive exploration of a long-horizon combinatorial solution space, taking into account the geometric relationships among each object. Furthermore, the placement locations in continuous space must also be considered to minimize end-effector displacement. Choosing the right placement of blocks can have a substantial impact on the later end-effector movement. The complexity of this task escalates as the number of objects involved increases.
b) Tower Construction X (Tower-X): X blocks are placed on the table, either in a stacked or unstacked configuration. The goal is to construct a tower following a specific order. Similar to the Pick-X task, the blocks may have varying heights, which means that a shorter block cannot be picked directly if a taller one is adjacent to it. Additionally, as mentioned in [11], both motion-level constraints and task-level constraints should be taken into account. For instance, if block A is initially atop block B, and the target order requires maintaining this arrangement at the target point, then A must be temporarily positioned at an intermediate point to first move block B and then A. Furthermore, in contrast to [11], we also consider the overall path optimality. This means that the placement of block B should be considered in the context of the entire trajectory to ensure minimum energy cost.
c) Tower Construction X with Tool (Tower-Tool-X): The blocks will be placed on the table within a large area, sometimes beyond the robot’s reach. Consequently, the robot must determine whether to use a tool (in the form of a hook) to pull objects into its reachable area before grasping them. This task is an expansion of the Tower-X challenge, incorporating a pulling action in addition to the standard grasping action. It also expands the workspace of a single robot arm to accomplish more complicated tasks. Furthermore, the robot must determine the target of the pulling phase to obtain a full optimal trajectory. We also demonstrate this task in real robot experiments.
@@ IV-B DTS vs MBTS
In this section, we compare our Dynamic Tree Search (DTS) with the Multi-Bound Tree Search (MBTS), which is widely used in previous LGP solvers. Our DTS adopts a similar idea of pose bound $\mathcal{P}_{pose}$ for quickly evaluating the feasibility of the current node but utilizes backward search from the target configuration. Therefore, we use $\mathcal{P}_{pose}$ for both MBTS and DTS to ensure a fair comparison. Moreover, considering that the parameter used to balance exploitation and exploration in MCTS/MBTS can significantly impact the results, we compare our DTS approach against three variations of MBTS. Each variant employs a distinct exploration-exploitation strategy, including greedy (MBTS-0), medium (MBTS-1) and lazy (MBTS-2). The comparison results are presented in Table I. We find that our proposed method visits the fewest nodes in the shortest time and achieves a $100\%$ success rate for all tasks. Moreover, thanks to the backpropagation strategy, it is applicable to tasks with exceedingly long horizons.
@@ IV-C MIP vs SLSQP and IPOPT
To obtain these results, we employ DTS as the high-level task planner and subsequently utilize MIQP, IPOPT, and SLSQP for low-level motion planning on the benchmarks with an increasing number of blocks. The implementation is based on Gurobi [16], Casadi [19], and Scipy [20], respectively. The objective of motion planning here is to determine the placement locations of the picked object. This task is particularly challenging due to the non-convex nature of the set of feasible points outside of a closed, bounded obstacle. As more obstacles are introduced, the problems become increasingly complex.
Successful planning is defined as obtaining a feasible solution within 10 seconds. We can observe from Table II that for setups with a small number of objects, all three solvers can easily achieve a high success rate within a short computation time. However, the success rates of the two NLP solvers decrease dramatically when more blocks are involved, indicating that NLP solvers struggle with complex non-convex problems. In contrast, MIQP can still provide optimal results. This highlights the fact that, given a feasible action skeleton, NLP solvers may become trapped in poor local optima or even fail to obtain a feasible solution. Consequently, the high-level task planner receives incorrect guidance, leading to a failure to identify the genuinely feasible sequence.
These findings highlight our efforts to introduce global optimization into the LGP framework, facilitating better communication between the high-level task planner and the low-level motion planner.
@@ IV-D D-LGP Performance
We conducted a detailed performance analysis of our proposed algorithm on various benchmarks, and the results are summarized in Table III. Our findings indicate that using $\mathcal{P}_{pose}$ for DTS significantly accelerates the hybrid programming, resulting in a feasible action skeleton and suboptimal motion variables. Given that TAMP involves long-horizon planning to achieve the lowest cost along the entire path, we subsequently perform full optimization based on the generated task sequence. When comparing the end-effector displacement between DTS and full optimization levels, we observed that full optimization can indeed yield an optimal full trajectory, whereas DTS assumes prior movements to be perfect and aims to rapidly find the optimal solution for the current step. However, it is worth noting that full optimization requires more computational time. This is primarily due to the exponential increase in binary variables, which poses challenges for modern MIQP solvers like Gurobi. With a greater number of objects, full optimization may encounter difficulties. This can be solved in the future by employing Iterative Regional Inflation by Semidefinite programming (IRIS) [21] to simplify MIQP formulation. Additionally, we provide the makespan of the plans, which closely aligns with the number of objects involved. This demonstrates that our generated plans prioritize executing actions that are essential rather than including feasible but extraneous actions.
@@ IV-E Real robot experiments
As mentioned in Section III-C, the output of D-LGP is defined in the object-centric frame, enabling adaptive low-level adjustments. For example, if the location of the target block changes, the robot’s end-effector can continue tracking it until picking it up. Moreover, since D-LGP runs significantly faster compared to standard LGP solvers, we have formulated it in a receding-horizon planning fashion. The D-LGP solver runs at approximately 10Hz, generating long-horizon task and motion sequences. We apply the first output to the robot while continuing to run D-LGP with the new state obtained from the 30Hz camera. This framework enables high-level task replanning in response to inaccurate execution (e.g. if a block falls) and external disturbances (e.g., random shifts in stacked block orders). We demonstrate these behaviors in our accompanying video.
@@ V Conclusion and Future Work
In this paper, we introduce a novel LGP framework called D-LGP, which leverages Dynamic Tree Search (DTS) and global optimization for efficient TAMP. We demonstrate that DTS enables target-oriented search and eliminates constraints on horizon length, allowing for the handling of tasks with extended horizons. We also show that our integrated global optimization formulation is capable of quickly obtaining global optima if the problem is feasible, evaluating the output of DTS effectively. We tested this approach on various benchmarks, highlighting its strong performance compared to state-of-the-art methods.
In this work, we used separate integer variables to represent each subspace corresponding to every face of the object. The number of integer variables grows exponentially with the number of objects, making it impractical to solve. This issue could be mitigated by utilizing a single integer variable to represent each convex region in the obstacle-free space.
Currently, our framework is tailored for tasks with explicit target configurations and does not support symbolic target descriptions, like “build the tower as high as possible”. However, we believe that our proposed method can be adapted for such cases by using sampling-based methods to determine the target configuration first.

### v3

@@ III-C D-LGP Algorithm
To tackle (1) iteratively from the target configuration, we introduce D-LGP in Alg. 2. Here, we employ a pseudoinverse successor function, denoted as $succ^{\dagger}(s_{k},s_{k+1})$, to deduce the goal-oriented action $a_{k+1}$ for transitioning from $s_{k}$ to $s_{k+1}$. This function disregards the symbolic and geometric constraints, focusing solely on determining the necessary action to reach the target state $s_{k+1}$ from $s_{k}$.
@@
2:$\bm{x}(kT):$ current geometric state, $s_{k}:$ current symbolic state, $g_{K}:$ target configuration
@@
5:$s_{K}\in\mathcal{G}(g_{K})$
6:$a_{k+1}=succ^{\dagger}(s_{k},s_{K})$
7:$\mathcal{B}_{a}=[a_{k+1}],\mathcal{B}_{g}=[g_{K}]$
8:while $(\bm{x}(kT),a_{k+1})$ infeasible do
9:   $g^{*}=sub\_goal(\bm{x}(kT),a_{k+1})$triangleright subgoal reasoning
10:   $s^{*}\in\mathcal{G}(g^{*})$
11:   $a_{k+1}=succ^{\dagger}(s_{k},s^{*})$
12:   $\mathcal{B}_{a}.append(a_{k+1})$
13:   $\mathcal{B}_{g}.append(g^{*})$
14:end while
@@
1:
2:$\bm{x}_{0}:$ initial configuration, $s_{0}:$ initial symbolic state, $g_{K}:$ target configuration
3:$a_{1:K}:$ action skeleton, $\bm{x}(t):$ full path
4:         Level 1: Dynamic Tree Search
5:Initialize:
@@
7:$\mathcal{B}_{a},\mathcal{B}_{g}\leftarrow\text{TaskGraph}(\bm{x}_{0},s_{0},g_{K})$
8:while not solved do
9:   $a_{k}=\mathcal{B}_{a}[-1]$ (feasible action given $\bm{x}_{k-1}$)
10:   $s_{k}=succ(s_{k-1},a_{k})$
11:   $g_{k}=\mathcal{B}_{g}[-1]$
12:   $\bm{x}(kT),\bm{u}(kT)=\text{MIQP}(a_{k},g_{k})$triangleright Eq. (4)
13:   $\mathcal{B}_{a},\mathcal{B}_{g}=\text{TaskGraph}(\bm{x}(kT),s_{k},g_{K})$triangleright see Alg. 1
14:   $k\leftarrow k+1$
15:end while
16:Level 2: Full Optimization
17:Given $a_{1:K}$:
18:
@@
20:
Given the current configuration $\bm{x}((k-1)T)$, we employ the conflict-driven TaskGraph algorithm (Alg. 1) to assess the feasibility of the proposed action. If the action is found infeasible, we generate a sequence of actions ($\mathcal{B}_{a}$) and subgoals ($\mathcal{B}_{g}$) to construct a tree connecting $s_{k-1}$ to $s_{K}$. $sub\_goal(\cdot)$ identifies the crucial subgoals that must be accomplished first to render action $a_{k}$ feasible, given the current configuration $\bm{x}((k-1)T)$. Subsequently, we calculate the continuous path from $t=(k-1)T$ to $t=kT$ based on (4). Since computing (4) is expensive, we decide to only focus on the final pose at $t=kT$ for fast feasibility checking and cost computation. Based on this, we can quickly solve it and obtain the placement locations $\bm{u}(kT)$ or returns global infeasibility.
(a) Initial and target states
(b) Solution
Figure 2: Illustrative example of Dynamic Tree Search. Given the initial and target configurations, $succ^{\dagger}(s_{0},s_{K})$ will infer the essential action as Pick[C #P3] (shown as dotted line in Fig. 2b). However, this is infeasible because B is on top of C. $sub\_{goal}(\cdot)$ will identify the essential subgoal as moving B to #p2 (given by MIQP) first. Then, C can be picked and placed at #p3, followed by picking B and A until reaching the final target.
Fig. 2 presents a simplified example to illustrate the running process of Dynamic Tree Search. Given the initial and target configurations, the essential action to connect them is Pick[C #P3], inferred through $succ^{\dagger}(s_{0},s_{K})$ that disregards all the constraints. It is clear that Pick[C #P3] is infeasible because B is on top of C. $sub\_{goal}(\cdot)$ will then be used to infer the appropriate configuration $g^{*}$ to make Pick C feasible. This process iterates until finding a feasible action respecting the current geometric configuration. In the provided example, the inferred action is Pick B, which is feasible concerning $\bm{x}(0)$. Then, Eq. (4) is invoked to generate the optimal placement #p2, ensuring minimal end-effector movement while satisfying geometric constraints. After executing the action, $s_{0}$ and $\bm{x}(0)$ update to $s_{1}$ and $\bm{x}(T)$, respectively. $succ^{\dagger}(\cdot,\cdot)$ is then used to determine the necessary action to achieve $s_{K}$ from $s_{1}$, resulting in Pick C, which is geometrically feasible w.r.t. $\bm{x}(T)$. The process continues similarly for Pick [B #p3] and Pick [A #p3], until reaching the final target configuration $g_{K}$.
After finding the full action skeleton $\{a_{1},\dots,a_{K}\}$, a full path optimization will be called to find the global optimal trajectory, focusing on the keyframes $t=0,T,\dots,KT$. This idea coincides with the pose bound $\mathcal{P}_{pose}$ and sequence bound $\mathcal{P}_{seq}$ used in multi-bound tree search [3], which proves using coarse path discretization can effectively speed up the optimization process. The difference is that we eliminate the full path bound $\mathcal{P}_{full}$ to fully utilize the robot controller to cope with uncertainties and external disturbance in the real world. Our TAMP approach runs in Cartesian space and relies on Operational Space Controller to actuate the robot joints, leading to reactive motion control. This is also highlighted in [18], where a Cartesian formulation of LGP was proposed.
@@ IV-A Benchmarks
We test our method in 3 different tabletop domains, all requiring long-horizon physical reasoning about the discrete actions and continuous motions. The reward is only given at the termination of the whole process. Across all task domains, we employ a random selection of 15 distinct initializations to assess the performance of our algorithm.
a) Obstructed Pick X (OP-X): In a cluttered environment, the robot’s task is to grasp an object surrounded by $X-1$ taller objects. If the taller objects are located near the target object, the robot must relocate them to create a clear path for grasping. This necessitates the robot’s comprehensive exploration of a long-horizon combinatorial solution space, taking into account the geometric relationships among each object. Furthermore, the placement locations in continuous space must also be considered to minimize end-effector displacement. Choosing the right placement of blocks can have a substantial impact on the later end-effector movement.
b) Tower Construction X (Tower-X): X blocks are placed on the table, either in a stacked or unstacked configuration. The goal is to construct a tower following a specific order. Similar to the Pick-X task, the blocks may have varying heights, which means that a shorter block cannot be picked directly if a taller one is adjacent to it. Additionally, as mentioned in [17], both motion-level constraints and task-level constraints should be taken into account. For instance, if block A is initially atop block B, and the target order requires maintaining this arrangement at the target point, then A must be temporarily positioned at an intermediate point to first move block B and then A. Furthermore, in contrast to [17], we also consider the overall path optimality. This means that the placement of block B should be considered in the entire trajectory to ensure minimum energy cost.
c) Tower Construction X with Tool (Tower-Tool-X): The blocks will be placed on the table within a large area, sometimes beyond the robot’s reach. Consequently, the robot must determine whether to use a tool (in the form of a hook) to pull objects into its reachable area before grasping them. This task is an expansion of the Tower-X scenario, incorporating a pulling action in addition to the standard grasping action. It also expands the workspace of a single robot arm to accomplish more complicated tasks. Furthermore, the robot must determine the target of the pulling phase to obtain a full optimal trajectory. We also demonstrate this task in real robot experiments.
@@ IV-B DTS vs MBTS
In this section, we compare our Dynamic Tree Search (DTS) with the Multi-Bound Tree Search (MBTS), which is widely used in previous LGP solvers. Our DTS adopts a similar idea of pose bound $\mathcal{P}_{pose}$ for quickly evaluating the feasibility of the current node but utilizes backward search from the target configuration. Therefore, we use $\mathcal{P}_{pose}$ for both MBTS and DTS to ensure a fair comparison. Moreover, considering that the parameter used to balance exploitation and exploration in MCTS/MBTS can significantly impact the results, we compare our DTS approach against three variations of MBTS. Each variant employs a distinct exploration-exploitation strategy, including greedy (MBTS-0), medium (MBTS-1) and lazy (MBTS-2). The comparison results are presented in Table I, with 100 seconds as the computation time threshold. We find that our proposed method visits the fewest nodes in the shortest time and achieves a $100\%$ success rate for all tasks. Moreover, thanks to the backpropagation strategy, it is applicable to tasks with exceedingly long horizons.
@@ IV-C MIP vs SLSQP and IPOPT
To obtain these results, we employ DTS as the high-level task planner and subsequently utilize MIQP, IPOPT, and SLSQP for low-level motion planning on the benchmarks with an increasing number of blocks. The implementation is based on Gurobi [15], CasADi [19], and Scipy [20], respectively. The objective of motion planning here is to determine the placement locations of the picked object. This task is particularly challenging due to the non-convex nature of the set of feasible solutions. As more obstacles are introduced, the problem becomes increasingly complex.
Successful planning is defined as obtaining a feasible solution within 10 seconds. We can observe from Table II that for setups with a small number of objects, all three solvers can easily achieve a high success rate within a short computation time. However, the success rates of the two NLP solvers decrease dramatically when more blocks are involved, indicating that NLP solvers struggle with complex non-convex problems. In contrast, MIQP can still provide optimal results. This highlights the fact that, given a feasible action skeleton, NLP solvers may be trapped in poor local optima or even fail to obtain a feasible solution. Consequently, the high-level task planner receives incorrect guidance, leading to a failure to identify the genuinely feasible sequence. Global optimization is therefore important to facilitate the communication between task planner and motion planner.
@@ IV-D D-LGP Performance
We conducted a detailed analysis of our proposed algorithm on various benchmarks. The results are summarized in Table III. Our findings indicate that using $\mathcal{P}_{pose}$ for DTS significantly accelerates the hybrid programming, resulting in a feasible action skeleton and suboptimal motion variables. Given that TAMP involves long-horizon planning to achieve the lowest cost along the entire path, we subsequently perform full optimization based on the generated task sequence. When comparing the end-effector displacement between DTS and full optimization levels, we observed that full optimization can indeed yield an optimal full trajectory, whereas DTS assumes prior movements to be perfect and aims to rapidly find the optimal solution for the current step. However, it is worth noting that full optimization requires more computational time. This is primarily due to the significant increase of binary variables, which poses challenges for modern MIQP solvers like Gurobi. With a greater number of objects, full optimization may encounter difficulties. This can be solved in the future by employing Iterative Regional Inflation by Semidefinite programming (IRIS) [21] to simplify MIQP formulation. Additionally, we provide the makespan of the plans, which closely aligns with the number of objects involved. This demonstrates that our generated plans prioritize executing actions that are essential rather than including feasible but extraneous actions.
@@ IV-E Real robot experiments
As mentioned in Section III-C, the output of D-LGP is defined in the object-centric frame, enabling adaptive low-level adjustments. For example, if the location of the target block changes, the robot’s end-effector can continue tracking it until picking it up. Moreover, since D-LGP runs significantly faster compared to standard LGP solvers, we have formulated it in closed-loop fashion. The D-LGP solver runs at approximately 10Hz, generating long-horizon task and motion sequences. We apply the first output to the robot while continuing to run D-LGP with the new state obtained from the 30Hz camera. This framework enables high-level task replanning in response to inaccurate execution (e.g., if a block falls) and external disturbances (e.g., random shifts in stacked block orders). We demonstrate these reactive behaviors in our accompanying video.
@@ V Conclusion and Future Work
In this paper, we introduce a novel LGP framework called D-LGP, which leverages Dynamic Tree Search (DTS) and global optimization for reactive TAMP. We demonstrate that DTS enables target-oriented search and eliminates constraints on horizon length, allowing for the handling of tasks with extended horizons. We also show that our integrated global optimization formulation is capable of quickly obtaining global optima if the problem is feasible, evaluating the output of DTS effectively. We tested this approach on various benchmarks, highlighting its strong performance compared to state-of-the-art methods.
In this work, we used separate integer variables to represent each subspace corresponding to every face of the object. The number of integer variables grows significantly with the number of objects, making it impractical to solve. This issue could be mitigated by utilizing a single integer variable to represent each convex region in the obstacle-free space.
Currently, our framework is tailored for tasks with explicit target configurations and does not support implicit target descriptions, like “build the tower as high as possible”. However, we believe that our method can be adapted for such cases by combining it with other approaches to determine the target configuration first.